OpenStudy (anonymous):

Does the equation represent a direct variation? If so, find the constant of variation. 5x = -3y

OpenStudy (anonymous):

Answers: A. no B. yes; k = -5/3 C. yes; k = 5/3 D. yes; k = 3/5

OpenStudy (anonymous):

Yes, can write as -3y = 5x y = (-5/3) *x the constant of variation is -5/3

OpenStudy (anonymous):

thank you so muh. this was the last question on my test!

OpenStudy (anonymous):

much*

OpenStudy (anonymous):

mhm :)